# Fourier123: One Image to High-Quality 3D Object Generation with Hybrid Fourier Score Distillation

### v1

@@ 4.1 Hybrid Fourier Score Distillation
where $\text{Re}(\mathbf{X})$ and $\text{Img}(\mathbf{X})$ denote the real and imaginary parts of $\mathbf{X}$ respectively. In this paper, the DFT is computed independently for each channel of feature maps.
Targeting at image-to-3D generation, we employ DFT to revisit the properties of the amplitude components (i.e., $\mathcal{A}(\mathbf{x})$), conducting detailed frequency analysis of images generated by 2D and 3D diffusion models. As shown in Fig. 2, the frequency amplitude of results from Zero123 is concentrated in low frequencies, while SD tends to produce higher frequency results, accompanied by better subjective quality. The main frequency discrepancy lies in the amplitude components. Meanwhile, in the spatial domain, the image generated by SD deviates from the input image, while Zero123 predicts a novel view that is geometrically more consistent with the input image.
Based on our observation, we propose employing a 3D diffusion model for supervision in the spatial domain, and a 2D diffusion model for supervising the amplitude component in the frequency domain. We focus on the amplitude component for two main reasons. Firstly, as previously mentioned, there is a significant disparity in the frequency amplitude results between 2D and 3D diffusion models. Secondly, the amplitude component represents the frequency distribution of the image, whereas the phase component more directly reflects the structure of the image. Since images produced by SD are visual-friendly but exhibit 3D inconsistent structure, we only adopt their amplitude components to learn their frequency distribution, which leads to fine details. We name this optimization function as 2D Fourier score distillation and formulate it as follows:
@@
where $\mathcal{A}(\cdot)$ indicates the amplitude component in the frequency domain. Intuitively, $\nabla_{\bm{\phi}}\mathcal{L}_{\textrm{2D-FSD}}$ converts the added noise $\bm{\epsilon}$ and predicted noise $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{y})$ into the Fourier space, and then calculates their amplitude components. For structural supervision, we utilize Zero123 in the spatial domain to construct 3D score distillation sampling, expressed as:
@@
As mentioned in Eq. 4, $\mathcal{C}(\mathbf{R},\mathbf{T})$ is the condition used in Zero123, which represents the camera pose of the generated novel view. By manipulating $\mathcal{C}(\mathbf{R},\mathbf{T})$, pseudo-GTs of different views can be generated for training, leading to structural constraints.
@@
Note that hy-FSD can be applied to any existing optimization-based generation method, replacing their score distillation functions in a plug-and-play manner. Experiments in Sec. 5.2 prove that hy-FSD brings significant performance gains to existing methods.
Figure 3: The workflow of Fourier123. We first use LGM [20] to initialize 3D Gaussian $\bm{\phi}$. Then, Zero123 [13] is used to supervise geometry in the spatial domain, while SD [41] supervises appearance in the frequency domain. The whole generation process takes less than one minute.
@@ 4.2 Overall Pipeline
Our Fourier123 pipeline is simple yet effective, which generates high-quality 3D assets based on a single image within 1 minute. It aims to combine the fast generation speed of large reconstruction models with the high-quality generation ability of optimization-based methods. Specifically, we employ 3D Gaussian as our 3D representation due to its superior optimization speed. Our overall framework consists of two parts: initialization with a large 3D reconstruction model and optimization through hy-FSD, as shown in Fig. 3. We parameterize the 3D Gaussians as $\bm{\phi}$. First, to initialize $\bm{\phi}$, we use the pre-trained large Gaussian model LGM [20], formulated as $F_{lgm}$, to generate coarse 3D Gaussian points based on the input image $y$, denoted as $\bm{\phi}=F_{lgm}(y)$. This stage is completed within 5 seconds. Then, we use Stable Diffusion (SD) [41] and Zero-1-to-3XL (Zero123) [13] to further optimize $\bm{\phi}$ via the proposed hy-FSD, obtaining the final result. This process takes about 52 seconds on a single NVIDIA 4090 GPU.
Gaussian initialization with large reconstruction model. Previous work [12; 18; 52] has proven that 3D initialization is critical to generation quality. As shown in Fig. 3(i), since we employ 3D Gaussian as the representation technique, the recently developed LGM is utilized to initialize a set of coarse Gaussians based on the input image, parameterized as $\bm{\phi}$. In this way, Fourier123 benefits from the advancements of large reconstruction models, accelerating 3D generation.
@@
Optimization with the hy-FSD. Although the initialization stage can generate coarse 3D assets with roughly reasonable geometry, their visual quality is unsatisfactory. To obtain high-quality 3D assets, we use the proposed hy-FSD to optimize $\bm{\phi}$. As depicted in the first term of Eq. 11, the 2D diffusion model (SD) is employed for appearance guidance. Note that since the classifier-free guidance [53] in SD is critical to its generation ability, a textual prompt is essential. This prompt can be generated by ChatGPT based on the input image, or it can be a universal text such as “A high-quality image”. The related workflow is shown in Fig. 3. We first render an image from $\bm{\phi}$, then add some noise and input it to SD. SD performs a denoising process for a few steps to get 2D supervision. Both the addition and removal of noise follow the DDIM schedule [40]. Considering that the products of SD tend to exhibit distorted but high-quality appearances, we extract their amplitude components in the frequency domain to utilize desired appearance priors while avoiding training being misled by the distorted content. This loss indirectly supervises 3D assets from the frequency perspective, avoiding conflicts between the generation domain of SD and other diffusion models or $\bm{\phi}$ itself. Benefiting from this, we can combine the 3D diffusion model (Zero123) more effectively.
@@
As depicted in the second term of Eq. 11, we use Zero123 to calculate 3D-SDS in the spatial domain, following similar operations as described in [13; 19]. Zero123 can generate novel views that are geometrically consistent with the input image, while strictly adhering to the given input camera pose. Taking advantage of this, $\bm{\phi}$ is trained with rich geometric priors and identity preservation.
@@
Figure 4: Visual results of ablation study on hy-FSD. We input a single image and a prompt, where the prompt is generated by ChatGPT based on the image. One can see that settings that use 2D-SDS to supervise in spatial domain all exhibit distortion. Our hy-FSD achieves the best results.
@@ 5 Experiment
In this section, we first introduce our implementation details, including parameter settings and evaluation metrics. Then we apply hy-FSD to existing well-known 3D generation pipelines, i.e., DreamFusion [7] and Magic123 [19], to affirm its universal efficacy. Finally, we evaluate the generation ability and runtime of the proposed Fourier123 compared with other SOTA methods.
@@ 5.1 Implementation Details
Optimizing the pipeline. We use the same settings for all objects. We only optimize the 3D Gaussian ellipsoids for 400 iterations, during which the learning rate of position information decays from $1\times 10^{-3}$ to $2\times 10^{-5}$. The Stable Diffusion [6] (SD) model of V2 is selected as 2D diffusion model. We set the scale for its classifier-free guidance to 7.5 because, on the one hand, a too high guidance scale (e.g., 100 used in [7] and [19]) leads to low generation quality [10]. On the other hand, in our method, the views generated by SD are used for supervision in the frequency domain rather than the RGB domain, which means guidance scale is not required to be too high for cross-view consistency at the pixel-level. Additionally, we leverage Zero-1-to-3XL for 3D structural supervision and its guidance scale is set to 5 according to [13]. Finally, the rendering resolution is set to $512\times 512$.
Camera setting. Due to the lack of pose onformation in the input reference image, we set its camera parameters as follows. First, we regard the reference image as being shot from the front view, i.e., azimuth angle is $0^{\circ}$ and polar angle is $90^{\circ}$, which is similar to [19]. Second, the camera is placed 1.5 meters from the content in the image, consistent with LGM [20]. Third, as we use Zero123 to supervise distillation, the field of view (FOV) of the camera is $49.1^{\circ}$.
Evaluation metrics and datasets. Following previous work [7; 19], we report CLIP-similarity [50] as the objective metric to measure the semantic distance between rendered images and input image. Besides, we collect two types of user scores from 40 volunteers as subjective metrics, which focus on the assessment of two critical aspects in the context of image-to-3D: reference view consistency (“User-Cons”) and overall generation quality (“User-Qual”). Both are rated from 1 (worst) to 5 (best). Following [7] and [19], we conduct ablation and comparison experiments on a dataset containing 51 images, which are collected from the GSO [54], OmniObject3D [55], and Internet.
Table 1: Quantitative results of ablation study, which are measured by CLIP-similarity $\uparrow$. The best and the second best results are highlighted in red and blue respectively.
@@
DreamFusion 0.6316 0.6213 0.6441 0.6312 0.6922
DreamGaussian 0.6961 0.7488 0.7634 0.6539 0.7927
@@
Fourier123 (Ours) Optimization-based 0.8010 4.5251 3.8333 57
5.2 Ablation on hy-FSD
To demonstrate that hy-FSD is an objective function with universal benefits for optimization-based generation methods, in this section, we apply hy-FSD to representative methods and analyze the impact of different score distillation functions on generation quality. DreamFusion (DF) [7] and DreamGaussian (DG) [11] are chosen since they have inspired numerous subsequent works and represent two optimization-based generation methods using NeRF and 3DGS as 3D representations. Applying hy-FSD to these two classic methods effectively demonstrates its broad applicability.
Both DF and DG utilize SDS for 3D generation. The difference is that DF only uses SD to provide 2D priors, while DG relies solely on Zero123 for image-to-3D task. To comprehensively validate the impact of score loss functions on generation quality and the effectiveness of the proposed hy-FSD, we conduct ablation experiments with five settings: “2D-SDS”: Vanilla DF has used 2D-SDS only. For DG, we substitute Zero123 with SD. “3D-SDS”: Vanilla DG has used 3D-SDS only. For DF, we replace SD with Zero123. “2D-SDS & 3D-SDS”: Following [19], we extend DF and DG to jointly use SD and Zero123 in spatial domain. “2D-SDS & 3D-FSD”: We use 2D priors of SD in the spatial domain and utilize 3D priors of Zero123 in the frequency domain. “2D-FSD & 3D-SDS” (Our hy-FSD): Distilling using SD in the frequency domain and Zero123 in the spatial domain.
We display visual results in Fig. 4, which reveal that using only 2D-SDS in the spatial domain cannot faithfully generate 3D assets, while adding 3D-SDS (2D-SDS & 3D-SDS) or using only 3D-SDS fails to produce high-quality objects. In addition, the setting that uses 2D-SDS and supervises with 3D diffusion model in the frequency domain (2D-SDS & 3D-FSD) results in a distorted structure. This proves that using 3D diffusion model in the frequency domain cannot effectively guide structure generation. Obviously, our hy-FSD setting, i.e., 2D-FSD & 3D-SDS, produces the best results with both high-quality appearances and natural geometry. We believe this is because the geometry is constrained at the pixel-level by 3D priors, while appearance features are supervised indirectly in the frequency domain by 2D diffusion model. We provide quantitative analysis using CLIP-similarity in Tab. 1 for a more objective illustration. An interesting finding is that 2D-SDS & 3D-SDS performs better than 3D-SDS, which proves that directly combining 2D and 3D diffusion models in the spatial domain can bring certain performance gains, consistent with the study of Magic123 [19]. However, the proposed hy-FSD brings the most noteworthy performance gains to existing pipelines, proving its effectiveness. We also conduct an ablation study on Fourier123 and provide results in the appendix.
Figure 5: Visual comparison. The input images are given on the left and runtime is listed below. Our method achieves better generation quality with competitive efficiency.
@@ 5.3 Comparison
As said in Sec. 4.2, text generated by ChatGPT and a universal text, “A high-quality image”, can both be used in SD. In this section, we display our results generated with universal text to prove that we can achieve SOTA effects without elaborated texts. However, using text produced by ChatGPT in SD actually performs better. We also showcase results of this optimal setting in Sec. D of appendix.
@@
Qualitative comparison. We provide qualitative comparisons of generation quality in Fig. 5. We primarily compare with three baselines from inference-only methods [20; 46; 47] and three optimization-based methods [11; 13; 19]. Rendering results from six samples are displayed for intuitive illustration. In terms of generation speed, our approach exhibits significant acceleration when compared to optimization-based methods. We attribute it to the initialization with a large reconstruction model and the convergence speed of 3D Gaussian. Regarding the quality of generated 3D assets, our method outperforms both inference-only and optimization-based methods, especially with respect to the fidelity of 3D geometry and visual appearance. In general, Fourier123 achieves a SOTA balance between generation quality and speed.
@@
Quantitative comparison. In Tab. 2, we report “CLIP-similarity” and two types of user scores to measure the generation ability of different methods. The average runtime on input with the resolution of $512\times 512$ is provided on the right to evaluate their efficiency. Note that the “Zero-1-to-3” in Tab. 2 refers to training a NeRF [22] with the pseudo-GTs from different views generated by Zero-1-to-3, which shares the same training settings as in [13]. One can see that our method outperforms the compared methods in terms of generation quality and achieves SOTA speed among optimization-based methods.
6 Conclusion and Limitation

### v2

@@ 4.1.1 Fourier Score for Stable Diffusion
where $\text{Re}(\mathbf{X})$ and $\text{Img}(\mathbf{X})$ denote the real and imaginary parts of $\mathbf{X}$ respectively.
Targeting at image-to-3D generation, we employ DFT to revisit the properties of the amplitude components (i.e., $\mathcal{A}(\mathbf{x})$), conducting frequency analysis of images generated by T2I and novel view diffusion models. As shown in Fig. 2, the frequency amplitude of results from Zero123 is concentrated in low frequencies, while SD tends to produce higher frequency results, accompanied by better subjective quality with finer details. Their discrepancy mainly lies in the middle amplitude.
Based on above observation, we employ the amplitude component of the T2I model for finer details. We focus on the frequency amplitude for two main reasons. 1) Phase component is related to the content structure of the image and amplitude component means texture features. The images generated by SD own impressive visual quality and fine details, but their structures are deviated from the input image. 2) Novel view diffusion model cannot provide detailed supervision since it only generates images corresponding to the input image, if input image is low-quality, Zero123 cannot improve visual quality, but SD can produce finer results. To improve quality during optimization, we have to choose amplitude component of T2I model for quality improvement. We name this optimization design as 2D Fourier score distillation and formulate it as follows:
@@
where $\mathcal{A}(\cdot)$ indicates the amplitude component in the frequency domain. Intuitively, $\nabla_{\bm{\phi}}\mathcal{L}_{\textrm{2D-FSD}}$ converts the added noise $\bm{\epsilon}$ and predicted noise $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t,\mathbf{y})$ into the frequency domain and calculates their differences of amplitude components, which is used to optimize 3D Gaussians $\bm{\phi}$.
4.1.2 Hybrid Fourier Score
Although $\nabla_{\bm{\phi}}\mathcal{L}_{\textrm{2D-FSD}}$ enables visual quality improvement during optimization, ensuring the generated 3D assets to be consistent with the input image is also essential. In addition, prolific structural supervision is also critical for 3D generation. To this end, we incorporate Zero123 into our distillation score. Specifically, we additionally utilize Zero123 in the spatial domain to construct 3D structure distillation sampling, expressed as:
@@ 4.1.2 Hybrid Fourier Score
As mentioned in Eq. 5, $\mathcal{C}(\mathbf{R},\mathbf{T})$ is the camera condition used in Zero123, which represents the camera pose of the generated novel view. By manipulating $\mathcal{C}(\mathbf{R},\mathbf{T})$, pseudo-GTs of different views can be generated for training, leading to structural constraints.
@@
Note that hy-FSD can be applied to any existing optimization-based generation method, replacing their score distillation functions in a plug-and-play manner. In Sec. 5.2, we conduct such experiments on NeRF [18] and 3DGS [29] respectively to prove our generalization and universality, proving that hy-FSD brings significant performance gains to existing methods.
Figure 3: The workflow of Fourier123. We first use $\mathcal{F}(\cdot)$ to initialize 3D Gaussian $\bm{\phi}$. $\mathcal{F}(\cdot)$ can be sphere initialization or large reconstruction model. Then, Zero123 [10] is used to supervise geometry in the spatial domain, while SD [8] supervises appearance in the frequency domain. The whole generation process takes less than one minute.
@@ 4.2 Overall Pipeline
Our Fourier123 pipeline is simple yet effective, which generates high-quality 3D assets based on a single image within 1 minute. We employ 3D Gaussian as our 3D representation due to its superior optimization speed. Our overall framework consists of two steps: initialization and optimization. As shown in Fig. 3, we parameterize the 3D Gaussians as $\bm{\phi}$. We first initialize $\bm{\phi}$ and this step is formulated as $\mathcal{F}(\cdot)$. $\mathcal{F}(\cdot)$ can be implemented by sphere initialization used in [34] or other point cloud generation models [13; 44]. The latter leads to better quality and is used in our main pipeline, but the former is still feasible. We use sphere initialization in the ablation experiments of Sec. 5.2. Next, we use Stable Diffusion (SD) [8] and Zero-1-to-3XL (Zero123) [10] to optimize $\bm{\phi}$ via the proposed hy-FSD, obtaining generated 3D assets. This process takes about 52 seconds on a single NVIDIA 4090 GPU.
Gaussian initialization. We employ differentiable 3D Gaussians as 3D representations. Their initialization is illustrated in Fig. 3 i), which can be modelled as: $\bm{\phi}=\mathcal{F}(\textbf{{I}})$, where I is the given input image. $\bm{\phi}$ is the initialized Gaussians, which will be iteratively optimized in subsequent processes (see Fig. 3 ii)). $\mathcal{F}$ is the initialization operation, and there are typically two choices: one is to randomly initialize Gaussians within a sphere in 3D space, as DreamGaussian [34] does. The other option is to employ a pre-trained large Gaussian generative model (e.g., LGM [44]) for initialization. In our implementation, we default to using LGM for initialization. However, our method is not initialization-sensitive, meaning that random sphere initialization is also feasible. To demonstrate the generalizability of our method, we analyze the initialization configurations in Sec. 5.2 and Sec. D of the Appendix.
@@
Optimization with the hy-FSD. Fourier123 uses hy-FSD to generate high-quality 3D assets with satisfactory appearances and geometries. Specifically, we use the proposed hy-FSD to optimize $\bm{\phi}$. As depicted in the first term of Eq. 11, SD is employed for appearance guidance with its generations in frequency amplitude. Note that since the classifier-free guidance [15] in SD is critical to its generation ability, a textual prompt is essential. This prompt can be generated by ChatGPT based on the input image, or it can be a universal text such as “A high-quality image”. We use the latter in this paper to prove superior convenience and generalization of our method, but in fact, using texts generated by ChatGPT leads to better performance and we conduct such experiments in Sec. F of our appendix. The detailed workflow is shown in Fig. 3. We first render an image from $\bm{\phi}$, then add some noise and input it to SD. SD performs a denoising process for a few steps to get 2D supervision. Both the addition and removal of noise follow the DDIM schedule [38]. Considering that the products of SD tend to exhibit distorted but high-quality appearances, we extract their amplitude components in the frequency domain to utilize desired appearance priors while avoiding training being misled by the distorted structure. This loss indirectly supervises 3D assets from the frequency perspective, avoiding conflicts between the generation priors of SD and other diffusion models or $\bm{\phi}$ itself. Benefiting from this, we can combine Zero123 more effectively.
@@
As depicted in the second term of Eq. 11, we use Zero123 to calculate 3D-SDS in the spatial domain, following similar operations as described in [10; 16]. Zero123 can generate novel views that are geometrically consistent with the input image, while strictly adhering to the given input camera pose. However, it cannot improve visual quality of the input image, but SD can. Taking advantage of hy-FSD, $\bm{\phi}$ is trained with rich geometric and appearance priors.
@@
Figure 4: Visual results of ablation study on hy-FSD. We input a single image and a prompt, where the prompt is generated by ChatGPT based on the image. One can see that settings that use 2D-SDS to supervise in spatial domain all exhibit content distortion. Our hy-FSD achieves the best results.
@@ 5.1 Implementation Details
Optimization details. We only optimize the 3D Gaussian ellipsoids for 400 iterations, where the learning rate of position information decays from $1\times 10^{-3}$ to $2\times 10^{-5}$. The Stable Diffusion [8] (SD) model of V2 is selected. We set its classifier-free guidance as 7.5 because, on the one hand, a too high guidance scale (e.g., 100 used in [14]) leads to low generation quality [7]. On the other hand, in our method, the views generated by SD are used for supervision in the frequency domain rather than the RGB domain, which means guidance scale is not required to be too high for cross-view consistency at the pixel-level. Additionally, we leverage Zero-1-to-3XL for 3D structural supervision and its guidance scale is set to 5 according to [10]. Finally, the rendering resolution is set to $512\times 512$.
Camera setting. Due to the lack of pose information in the input reference image, we set its camera parameters as follows. First, we regard the reference image as being shot from the front view, i.e., azimuth angle is $0^{\circ}$ and polar angle is $90^{\circ}$, which is similar to [16]. Second, the camera is placed 1.5 meters from the content in the image, consistent with LGM [44]. Third, as we use Zero123 to supervise distillation, the field of view (FOV) of the camera is $49.1^{\circ}$.
Evaluation metrics and datasets. Following previous work [14; 16], we report CLIP-similarity [48] as the objective metric to measure the semantic distance between rendered images and input image. Besides, we collect two types of user scores from 40 volunteers as subjective metrics, which focus on the assessment of two critical aspects in the context of image-to-3D: reference view consistency (“User-Cons”) and overall generation quality (“User-Qual”). Both are rated from 1 (worst) to 5 (best). Following [14] and [16], we conduct ablation and comparison experiments on a dataset containing 51 images, which are collected from the Objaverse [50], OmniObject3D [51], and Internet. Moreover, 100 3D objects are randomly selected from GSO [52] to evaluate performance with lateral Ground Truth, sothat we can provide image-level metrics for more objective comparison, including PSNR, SSIM [53] and LPIPS [54].
Table 1: Quantitative results of score function ablation, which are measured by CLIP-similarity $\uparrow$. The best and the second best results are highlighted in red and blue respectively.
@@
DreamFusion 0.6139 0.6950 0.6661 0.6533 0.7416
DreamGaussian 0.5612 0.6386 0.5923 0.5905 0.7546
5.2 Applying hy-FSD to Existing Methods
To demonstrate that hy-FSD is generally beneficial to optimization-based generation methods, in this section, we apply hy-FSD to representative methods and analyze the impact of different score distillation functions on generation quality. DreamFusion (DF) [14] and DreamGaussian (DG) [34] are chosen since they have inspired numerous subsequent works and represent two optimization-based generation methods using NeRF and 3DGS as 3D representations. Applying hy-FSD to these two classic methods effectively demonstrates its broad applicability.
Both DF and DG utilize SDS for 3D generation. The difference is that DF only uses SD to provide 2D priors, while DG relies solely on Zero123 for image-to-3D task. To comprehensively validate the impact of score loss functions on generation quality and the effectiveness of the proposed hy-FSD, we conduct ablation experiments with five settings: (a)“2D-SDS”: Vanilla DF has used 2D-SDS only. For DG, we substitute Zero123 with SD. (b)“3D-SDS”: Vanilla DG has used 3D-SDS only. For DF, we replace SD with Zero123. (c)“2D-SDS & 3D-SDS”: Following [16], we extend DF and DG to jointly use SD and Zero123 in spatial domain. (d)“2D-SDS & 3D-FSD”: We use 2D priors of SD in spatial domain and utilize 3D priors of Zero123 in frequency domain. (e)“2D-FSD & 3D-SDS” (hy-FSD): Distilling with SD in the frequency domain and Zero123 in the spatial domain.
We display visual results in Fig. 4. From the comparison between (a) and (b), one can see that (a) produces much worse structure compared with (b), which is consistent with the aforementioned statement that Zero123 can generate more suitable geometry than SD. Moreover, the structures of (c) are also worse than that of (b), such as the back views of the second and third columns. (c) even exhibits Janus problem in fox of DF, that is, producing two heads of the fox in the back view. This is because the 3D priors of Zero123 is corrupted by SD. If using SD in spatial domain but using Zero123 in frequency domain, (d) shows that 3D geometry even becomes worse. Compared with (c), the dinosaur and fox of (d) contain more distorted structures. In contrast, our method that uses SD in frequency domain and Zero123 in spatial domain, unleashes the generation ability of SD while benefiting from structure priors of Zero123. (e) performs the best visual quality than other ablation settings in both texture and structure levels.
We provide quantitative analysis using CLIP-similarity in Tab. 1 for a more objective illustration, average scores are calculated on 151 3D objects, including the collected 51 cases and 100 samples from GSO data. One can see that directly using zero123 can achieve acceptable image-to-3D effect, but with hy-FSD, we further bring notable performance gains to existing pipelines.
@@ 5.2 Applying hy-FSD to Existing Methods
Fourier123 (Ours) Optimization-based 0.8010 4.5251 3.8333 52
Table 3: More quantitative comparison results with lateral Ground Truth. Experiments are conducted on the GSO subset containing 100 random selected 3D objects. The best and the second best results are highlighted in red and blue respectively.
Methods LGM CRM InstantMesh Zero-1-to-3 Magic123 DreamGaussian Fourier123 (Ours)
PSNR $\uparrow$ 17.2192 16.8246 14.8517 14.8944 14.5827 16.4898 21.5049
SSIM $\uparrow$ 0.8342 0.8423 0.7942 0.8314 0.7704 0.8231 0.8650
LPIPS $\downarrow$ 0.1806 0.1853 0.2386 0.1832 0.2764 0.2113 0.1112
Figure 5: Visual comparison. Input images are given on the left and runtime is listed below. For clear comparison, we omit LGM and InstantMesh here and the full version can be found in Sec. A.
@@ 5.3 Comparison
As said in Sec. 4.2, text generated by ChatGPT and a universal text, “A high-quality image”, can both be used in SD. In this section, we display our results generated with universal text to prove that we can achieve SOTA effects without elaborated texts. However, using text produced by ChatGPT in SD actually performs better. We also showcase results of this optimal setting in Sec. F of appendix.
@@
Qualitative comparison. We provide qualitative comparisons of generation quality in Fig. 5. We primarily compare with three baselines from inference-only methods [44; 45; 46] and three optimization-based methods [34; 10; 16]. In terms of generation speed, our approach exhibits significant acceleration compared to optimization-based methods. Regarding the quality of generated 3D assets, our method outperforms both inference-only and optimization-based methods, especially with respect to the fidelity of 3D geometry and visual appearance. Zero123 and DreamGaussian both only use 3D SDS and Magic123 uses 2D and 3D SDS. One can see that appearances of Zero123 and DreamGaussian tend to be smooth and blurry, which are influenced by the limited quality pseudo-GTs generated by Zero123. Magic123 produces relatively clear results, but its geometric structures are distorted, such as the first, second, and last rows. This is because results from SD corrupts the 3D priors of Zero123. In general, Fourier123 achieves SOTA balance between appearance and structure.
@@
Quantitative comparison. We first use the collected 51 3D objects to evaluate performance. In Tab. 2, we report “CLIP-similarity” and two types of user scores to measure the generation ability of different methods. The average runtime on input with the resolution of $512\times 512$ is provided on the right to evaluate their efficiency. Note that the “Zero-1-to-3” in Tab. 2 refers to training a NeRF [18] with the pseudo-GTs from different views generated by Zero-1-to-3, which shares the same training settings as in [10]. One can see that our method outperforms the compared methods in terms of generation quality and achieves SOTA speed among optimization-based methods.
To enable objective image-level evaluation with lateral Ground Truth, we further use 100 3D objects from GSO dataset. Considering that both Wonder3D [40] and CRM [45] selected 30 cases from GSO, we believe the subset we used is sufficient for quantitative comparison. PSNR, SSIM, and LPIPS are reported in Tab. 3. It is easy to see that our method performs superior 3D generation ability compared to existing baselines.
6 Conclusion